The typical​ (median) price of a​ single-family home in a certain region rose from ​$77 comma 600 in 1992 to ​$184 comma 000 in 2007. Calculate the relative change in price from 1992 to 2007​, and compare it to the overall rate of inflation as measured by the Consumer Price Index.
Answer #1

Solution:

Relative price change = ((Newest -Oldest)/Oldest)×100

= ((184-77.6)/77.6)×100

= 1.3711×100 = 137.11

Consumer price index can be calculated as

= (New Price/Old Price)×100

= (184/77.6)×100

= (2.3711)×100

= 237.11